geerky42 (geerky42):

How can I derive closed form of polynomial recursive sequence? Saying we have \[\huge a_n = a_{n-1} + n^2 + 4n,\text{ where }a_0=0\] What is its closed form?

OpenStudy (amistre64):

solve the homogenous, then solve for the particular ... difference equations are similar to differential equations

OpenStudy (amistre64):

or, define a few terms and see if we get to a constant difference tier

geerky42 (geerky42):

Well, I'm not familiar with differential equations yet. Can you show me exactly how to do second approach?

geerky42 (geerky42):

I'm self learning, so I may need to learn something beforehand, but I dunno.

OpenStudy (amistre64):

the most basic approach is to list your first few terms, say 5 up to 10? then see if we can develop the difference teirs

geerky42 (geerky42):

Basically like this?\[a_n = a_{n-1} + n^2+4n = a_{n-2}+(n-1)^2+4(n-1)+ n^2+4n=\cdots\] And sorry, but exactly what is "difference teirs?"

OpenStudy (amistre64):

the other way is to define the homogenous \[ a_n - a_{n-1} = 0\] and use it in determining the particular let an = x^n \[ x^n - x^{n-1} = 0\] \[ x^{n-1}( x - 1) = 0~:~x=1\] \[c_0(1)^{1}=c_1(1)^{0}+(1)^2+4(1)\] \[c_0(1)^{2}=c_1(1)^{1}+(2)^2+4(2)\] seems familiar

geerky42 (geerky42):

Um let's just stick to most basic approach because I'm lost already.

OpenStudy (amistre64):

well, lets define the terms a0, a1, a2, a3, a4, a5, a6, .. as many as you can muster

geerky42 (geerky42):

By define, you mean find the value of a0, a1, a2, a3, etc?

OpenStudy (amistre64):

yes 0 1 2 3 4 ... 0 5 17 38 70 ...

geerky42 (geerky42):

Ok thanks for doing calculations for me :) so what's next?

geerky42 (geerky42):

How many terms should I define up to?

OpenStudy (amistre64):

0 | 5 | 17 | 38 | 70 5 12 21 32 7 9 11 2 2 take the differences, and then the differences of the differecnes, and then .... continue doing this in teirs until you come to a constant tier, one where the differences are all the same

geerky42 (geerky42):

ah I see

OpenStudy (amistre64):

define a5 and see if we get to a constant of 2 11+2 = 13 +32 = 45 + 70 = 115 is a5 = 115?

OpenStudy (amistre64):

70 + 5^2 + 4(5) 70 + 45, it is :)

geerky42 (geerky42):

Yeah I checked

OpenStudy (amistre64):

now we can take the values on the front of each row and work it into a formula to find the(n-1)th term on the top

OpenStudy (amistre64):

one way is to do a formula, im curious if we can get it this way tho notice that if we were to represent this by a polnomial f(x) = a + bx + cx^2 + dx^3 + ex^4 + ... we have 3 tiers, so the thrid derivative is constant f'''(x) = 2 f''(x) = 2x + k f'(x) = x^2 + kx + j f(x) = 1/3 x^3 + 1/2 kx^2 + jx + m ------------------------------- f(0) = 0, so m=0 ------------------------ f(1) = 5 = 1/3 + 1/2 k + j ------------------------ f(2) = 17 = 1/3 (2^3) + 1/2 k (2^2) + 2j ----------------------- think thisll work? 2 equations in 2 unknowns?

geerky42 (geerky42):

Hmm interesting... Give me couple mins to solve for unknowns.

geerky42 (geerky42):

I got \(k=5\) and \(j=\dfrac{13}{6}\) So closed form would be \(f(n)=\boxed{a_n=\dfrac{1}{3}n^3+\dfrac{5}{2}n^2+\dfrac{13}{6}n}\) Checking: from closed form, \(a_6 = 175\); from recursive form, \(a_6 = 115 + 6^2 + 4(6) = 175\) It seems to work. Awesome!

geerky42 (geerky42):

I'm surprised to see that coefficients are ugly fraction, yet it will always result as integer for \(n\in\mathbb N\)

OpenStudy (amistre64):

yeah, i even impressed myself :) i know that differences are related to derivatives, so i took a stab at it

geerky42 (geerky42):

Awesome. Leveled up: math skill: +1 thanks to you! :)

OpenStudy (amistre64):

youre welcome :)